# MRGen: Diffusion-based Controllable Data Engine for MRI Segmentation towards Unannotated Modalities

### v1

@@ 5.2 Quantitative Results
Image generation. As shown in Table 2, images from the source domain exhibit high FID values with the target domain, indicating substantial discrepancies between distinct radiology modalities. Similarly, images augmented by DualNorm and translated by CycleGAN also exhibit high FID values, confirming their limited ability to emulate target-domain images. Conversely, our MRGen presents a significantly lower FID, demonstrating its capability to accurately generate images in target modalities, thus providing a foundation for training segmentation models with synthetic data.
Segmentation. As presented in Table 3, we have the following four key observations: (i) The significant discrepancies among different modalities hinder the ability of nnUNet and UMamba trained exclusively on source-domain data to generalize to the target domain, resulting in the lowest average DSC scores. (ii) The conventional image translation approach, CycleGAN, and the augmentation-based method, DualNorm, achieve moderate improvements in average DSC scores. However, these methods still struggle in the target domain and occasionally underperform in specific scenarios. (iii) In contrast, our proposed MRGen generates high-quality images for target domains, achieving the highest DSC score in 11 out of 13 experiments and significantly outperforming CycleGAN. Notably, MRGen consistently improves performance across all settings compared to source-only training, demonstrating the adaptability and effectiveness of its synthetic data across various segmentation architectures. (iv) The interactive SAM2 model, trained on substantially larger datasets, exhibits robust zero-shot segmentation capabilities with perturbed oracle box prompts. However, as a semi-automatic method, its reliance on spatial prompts and manual intervention limits its scalability and practical applicability.
@@ 5.3 Qualitative Results
Generation. As depicted in Figure 4, the images of distinct modalities exhibit substantial visual discrepancies, making it challenging for DualNorm to simulate via augmentation. Additionally, CycleGAN is prone to instability during training, frequently leading to model collapse when learning such complex translations. Conversely, our proposed MRGen effectively generates images that closely resemble the target domain and align with conditioned organ masks. This capability provides a solid foundation for the controlled generation of training data to extend segmentation models towards unannotated modalities.
Segmentation. As shown in Figure 5, despite significant variability between the source and target domains, MRGen substantially improves segmentation accuracy across all organs with high-quality synthetic data. However, the data derived from DualNorm and CycleGAN lead to unsatisfactory segmentation results, due to notable discrepancies from the target domain or low quality.
@@
Table 4: Ablation Study on Segmentation Performance (DSC score). Here, T2S. denotes T2-SPIR.
5.4 Ablation Studies
To validate the effectiveness of our modules and strategies, we conduct ablation studies on the downstream segmentation task. Specifically, we train nnUNet using data synthesized under different training and inference settings. As presented in Table 4, we have the following observations: (i) even without integrating unannotated target-domain data during training, MRGen still improves segmentation performance on the target modality. This indicates that the two-stage training strategy effectively enables MRGen to extend controllable generation to modalities without mask annotations; (ii) the autofilter pipeline based on SAM2 further improves performance by selecting data pairs that align better with the mask conditions. More ablation studies on image generation will be provided in Section D.2 of the Appendix.
@@ 6 Conclusion
This paper explores a novel paradigm of applying generative models in medical imaging: controllably synthesizing data for modalities lacking mask annotations, to train MRI segmentation models towards these challenging scenarios. To support this, we curate a large-scale radiology image-text dataset, MedGen-1M, featuring detailed modality labels, attributes, regions, and organ information, along with a subset of organ mask annotations. Built upon the dataset, our diffusion-based controllable data engine, MRGen, generates MR images conditioned on text prompts and masks, enabling data synthesis across unannotated modalities for training segmentation models. Comprehensive quantitative and qualitative evaluations across diverse modalities demonstrate that MRGen can effectively synthesize data, advancing MRI segmentation for unannotated modalities.
@@ Appendix A Preliminaries on Diffusion Models
Appendix B Details of MedGen-1M
In this section, we provide additional details about our collected and curated MedGen-1M dataset. In Section B.1, we elaborate on the implementation details of the automatic annotation pipeline; and in Section B.2, we present more comprehensive statistics about the dataset.
@@ B.1 Automatic Annotations
We employ an automated annotation pipeline to annotate the data in our MedGen-1M, ensuring that the samples contain comprehensive clinically relevant annotation information. This process primarily includes two components: human body region classification and modality explanation, which will be detailed as follows.
@@ B.2 Dataset Statistics
In this section, we present more detailed statistics about our curated MedGen-1M dataset, including the unannotated image-text pairs from Radiopaedia333radiopaeida.org, as well as the mask-annotated data sourced from various open-source datasets.
Data without mask annotations. For the image-text pairs from Radiopaedia-CT and Radiopaedia-MRI, which are used for training the autoencoder and text-guided generation, we allocate 1% of the data as a test set to evaluate reconstruction and generation performance, maximizing the amount of data available for pretraining. As a result, 1,007,616 samples are used for training, comprising 804,628 from Radiopaedia-CT and 202,988 from Radiopaedia-MRI. The test set consists of 10,179 samples, with 8,128 from Radiopaedia-CT and 2,051 from Radiopaedia-MRI.
We conduct a statistical analysis of the distribution of modalities in Radiopaedia-MRI, as presented in Figure 6 (a). The free-text modality labels cover approximately 300 categories, providing a diverse set of MRI modalities that form a crucial foundation for MRGen to learn text-guided generation and expand its mask-conditioned generation capabilities towards modalities originally lacking mask annotations. Furthermore, the distribution of images across different regions in Radiopaedia-CT and Radiopaedia-MRI is presented in Figure 6 (b).
Figure 6: Data Statistics of Radiopaedia. (a) Distribution of slice counts across various modalities in Radiopaedia-MRI; (b) Proportional distribution of slices across different regions in Radiopaedia-CT and Radiopaedia-MRI.
@@
MSD-Liver [2] Liver CT 105 46,695 15,260 26 11943 3903
AMOS22CT [23] Liver, Right Kidney, CT 200 26,069 19,023 100 15,361 11,256
Left Kidney, Spleen, etc.*
Total / / 1,083 105,074 48,501 218 34,907 18,802
Table 5: Details of Segmentation-annotated Datasets in MedGen-1M. Here, # Vol. represents the number of 3D Volumes, # Slc. denotes the number of 2D slices, and # Slc. w/ mask indicates the number of 2D slices with mask annotations. *: AMOS22CT [23] also contains images and mask annotations for the following organs: Gallbladder, Esophagus, Stomach, Aorta, Inferior Vena Cava, Pancreas, Duodenum, Urinary Bladder, and Adrenal Gland.
@@ Appendix C Implementation Details
In this section, we will provide a comprehensive explanation of the implementation details discussed in the paper. Specifically, Section C.1 describes the preprocessing and augmentation strategies applied to the training data. Section C.2 elaborates on the details of the autofilter pipeline. Finally, Section C.3 outlines the implementation details of the baselines.
@@ C.1 Preprocessing & Augmentation
Data preprocessing. To ensure consistency across data from various sources and modalities, we apply tailored preprocessing strategies as follows: (i) For data from Radiopaedia, the images are directly rescaled to the range [0, 1]; (ii) For CT images with mask annotations, intensities are clipped to [-300, 200] and then rescaled to [0, 1]; (iii) For MR images with mask annotations, intensities are clipped to the 0.5 and 99.5 percentiles and rescaled to [0, 1]. After normalization, all data are subsequently rescaled to [-1, 1] for training various components of MRGen, including the autoencoder, diffusion UNet, and mask condition controller. For training downstream segmentation models, images are rescaled to [0, 255] and saved in ‘.png’ format, followed by the preprocessing configurations of nnUNet [21] and UMamba [34].
Data augmentation. During autoencoder training, we apply random data augmentations to the images with a 20% probability. These augmentations included horizontal flipping, vertical flipping, and rotations of 90, 180, 270. In contrast, no data augmentations are applied during the training of the diffusion UNet and mask condition controller. For downstream segmentation models, we adhere to the default data augmentation strategies provided by nnUNet [21] and UMamba [34].
@@ C.2 Autofilter Pipeline
When deploying our proposed data engine, MRGen, to synthesize data for segmentation model training, we adopt the off-the-shelf SAM2 [41] model to perform automatic interactive segmentation on generated images, with the mask conditions as prompt. This pipeline automatically selects samples that are fidelity to the condition masks, thus ensuring the quality of synthetic image-mask pairs. In this section, we elaborate on more implementation details of this automatic filtering pipeline, particularly focusing on the generation of MR images that encompass multiple organs.
Specifically, we begin by defining the following thresholds: confidence threshold ($\tau_{\mathrm{conf}}$), IoU score threshold ($\tau_{\mathrm{IoU}}$), average confidence threshold ($\bar{\tau}_{\mathrm{conf}}$), and average IoU threshold ($\bar{\tau}_{\mathrm{IoU}}$). Both the mask ($\mathcal{M^{\prime}}_{t}$) and the generated image ($\mathcal{I^{\prime}}_{t}$) are fed into SAM2. For each organ mask $\mathcal{M^{\prime}}^{i}_{t}$ in $\mathcal{M^{\prime}}_{t}$, SAM2 will output a confidence score ($s^{i}_{\mathrm{conf}}$) and a pseudo mask annotation, which is then used to calculate the IoU score ($s^{i}_{\mathrm{IoU}}$) with $\mathcal{M^{\prime}}^{i}_{t}$. For each generated sample ($\mathcal{I^{\prime}}_{t}$), it is regarded as a high-quality sample if the following conditions are satisfied: $\{s^{i}_{\mathrm{IoU}}\geq\tau_{\mathrm{IoU}},s^{i}_{\mathrm{conf}}\geq\tau_{\mathrm{conf}}\,|\,\forall i\}$, and $\{\bar{s}_{\mathrm{IoU}}\geq\bar{\tau}_{\mathrm{IoU}}$, $\bar{s}_{\mathrm{conf}}\geq\bar{\tau}_{\mathrm{conf}}\}$. Otherwise, the sample will be discarded.
In all experiments, the predefined thresholds are set as follows: $\tau_{\mathrm{IoU}}=0.70$, $\tau_{\mathrm{conf}}=0.80$, $\bar{\tau}_{\mathrm{IoU}}=0.80$, and $\bar{\tau}_{\mathrm{conf}}=0.90$.
@@ C.3 Baselines
CycleGAN. We follow the official implementation and training strategies across all experimental settings. Subsequently, source-domain images are translated into the target domain and paired with the source-domain masks to create paired samples for training downstream segmentation models.
SAM2. We use SAM2-Large in all the experiments. Segmentation results are derived in a slice-by-slice and organ-by-organ manner: For each slice with mask annotations, we simulate box prompts for each annotated organ individually. To take into account the error introduced by manual intervention, the oracle boxes are randomly shifted at each corner, by up to 8% of the image resolution, following the protocol in MedSAM [33].
DualNorm. Following the official implementation, we apply random non-linear augmentation on each image in the source domain, to generate a source-dissimilar training sample, and train the dual-normalization model. All preprocessing steps, network architectures, and training strategies follow the official recommendations, with the exception that images are rescaled to 512 $\times$ 512, consistent with other approaches. Note that we evaluate DualNorm on all the slices in the test set, offering a more rigorous evaluation compared to the official implementation, which only considers slices with segmentation annotation.

### v2

@@ 4.2 Quantitative Results
Generation. As shown in Table 2, source-domain images exhibit high FID values compared to the target domain, indicating substantial discrepancies across distinct modalities. Similarly, images augmented by DualNorm and translated by CycleGAN and UNSB also show high FID values, confirming their limited ability to emulate target-domain images. In contrast, our MRGen presents a significantly lower FID, demonstrating its ability to accurately generate images of target modalities, providing a foundation for training segmentation models with high-quality synthetic data.
Segmentation. As depicted in Table 3, we draw the following three observations: (i) significant discrepancies between different modalities challenge the generalization ability of nnUNet and UMamba models trained solely on source-domain data, leading to notably lower average DSC scores; (ii) DualNorm and segmentation models trained with data translated by CycleGAN and UNSB, achieve slight or moderate improvements in average DSC scores, but consistently underperform across most scenarios; (iii) conversely, our MRGen produces high-quality target-domain samples for training segmentation models, thus achieving the highest DSC score in 8 out of 10 experiments, significantly outperforming others. Notably, MRGen consistently improves performance of nnUNet and UMamba, demonstrating the adaptability and robustness of its synthetic data across various segmentation architectures. More comparisons will be included in the Sec. D.2.
@@ 4.3 Qualitative Results
Generation. As presented in Figure 4, images of distinct modalities exhibit substantial visual discrepancies, making it challenging for DualNorm to simulate via exhaustive augmentation. While CycleGAN and UNSB preserve contours well, they suffer from unstable training and model collapse when learning complex transformations, thus failing to synthesize target-domain images accurately. In contrast, MRGen effectively generates images that closely resemble target domains and align with given organ masks, providing compelling evidence for controllable MRI data synthesis.
Figure 5: Qualitative Results of Segmentation towards Unannotated Modalities. Significant imaging differences between source-domain ($\mathcal{D}_{s}$) and target-domain ($\mathcal{D}_{t}$) make segmentation on target domains ($\mathcal{D}_{t}$) extremely challenging. We visualize liver, right kidney, left kidney and spleen in the first two rows, prostate in the third row, and pancreas in the fourth row with distinct colors.
Segmentation. As illustrated in Figure 5, despite significant appearance variations among distinct modalities, MRGen substantially improves segmentation accuracy across all organs with high-quality synthetic data. However, DualNorm, and segmentation models trained on data derived from CycleGAN and UNSB, yield unsatisfactory results.
Model SDM [49] SDM-ft MRGen-M MRGen
PSNR $\uparrow$ 31.32 35.65 42.62
SSIM $\uparrow$ 0.989 0.996 0.999
MSE $\downarrow$ 0.0037 0.0014 0.0003
FID $\downarrow$ 249.24 91.48 41.82 39.63
CLIP-I $\uparrow$ 0.3151 0.6698 0.7512 0.8457
CLIP-T $\uparrow$ 0.1748 0.3199 0.3765 0.3777
Table 4: Ablation on Reconstruction and Text-guided Generation. Here, MRGen-M adopts the same autoencoder as MRGen.
4.4 Ablation Studies
To validate the effectiveness of our strategies and modules, we conduct comprehensive ablation experiments on both generation and downstream segmentation task, as follows.
Generation. We assess autoencoder reconstruction and text-guided generation performance on MRGen-DB test set across various models, including: (i) pretrained Stable Diffusion (SDM), (ii) SDM finetuned on MRGen-DB (SDM-ft), (iii) our model conditioned only on free-text modality labels (MRGen-M), and (iv) our full MRGen with text prompts, incorporating modalities, attributes, regions, and organs information. The reconstruction quality is assessed using PSNR, SSIM, and MSE between the reconstructed and original images. For generation, we use FID, along with image-to-image (CLIP-I) and image-to-text (CLIP-T) similarities between generated images and ground truth or modality labels, computed by BiomedCLIP [62].
As presented in Table 4, we can draw the following observations: (i) while SDM pretrained on natural images performs poorly on MRI, finetuning on MRGen-DB yields substantial improvement in both reconstruction and generation; (ii) MRGen, with higher latent dimensions (16 versus 4 of SDM) and the BiomedCLIP text encoder, achieves significantly better performance; and (iii) our templated text prompts further enable MRGen to better distinguish distinct modalities, regions, and organs, leading to superior synthesis quality. To this end, we employ high-capacity autoencoders trained on MRGen-DB, text encoder pretrained on biomedical data, and clinically relevant templated text prompts to ensure accurate and realistic MRI synthesis.
@@ 4.4 Ablation Studies
Table 5: Ablation on Segmentation Performance (DSC score).
Segmentation. We train nnUNet with data generated by MRGen under various training and inference settings. As shown in Table 5, we have the following two observations: (i) MRGen boosts segmentation performance with synthetic data, even without incorporating target-domain unannotated data during training, demonstrating its strong generalization capability to underrepresented and annotation-scarce modalities; and (ii) the inclusion of target-domain unannotated images and the autofilter pipeline further improve performance by mitigating overfitting and selecting high-quality samples aligned with mask conditions.
@@ 5 Conclusion
This paper explores generative models for controllable MRI generation, particularly to facilitate training segmentation models for underrepresented modalities lacking mask annotations. To support this, we curate a large-scale radiology image-text dataset, MRGen-DB, featuring detailed modality labels, attributes, regions, and organ information, with a subset of organ mask annotations. Built on this, our diffusion-based data engine, MRGen, synthesizes MR images of various annotation-scarce modalities conditioned on text prompts and masks. Comprehensive evaluations across diverse modalities demonstrate that MRGen effectively improves segmentation performance on unannotated modalities by producing high-quality synthetic data. We believe this will offer new insights into addressing the scarcity of medical data and annotations, holding clinical significance.
@@ Appendix A Preliminaries on Diffusion Models
Appendix B Details of MRGen-DB & Synthetic Data
In this section, we provide additional details about our collected and curated MRGen-DB dataset. In Sec. B.1, we elaborate on the implementation details of the automatic annotation pipeline; and in Sec. B.2, we present more comprehensive dataset statistics. Additionally, in Sec. B.3, we provide statistics on the MRGen-synthesized data used for downstream segmentation models training under each experimental setting.
@@ B.1 Automatic Annotations
We employ an automated annotation pipeline to annotate our MRGen-DB dataset, ensuring that the templated text prompts contain sufficient and comprehensive clinically relevant information to distinguish distinct modalities, regions, and organs. This process primarily consists of two precise and controllable components: human body region classification and modality explanation, which will be detailed as follows.
@@
To ensure reliability and accuracy, we have randomly and uniformly sampled approximately 2% (5K out of 250K) of region annotations and 20% (60 out of $\sim$300) of modality attribute annotations for manual verification, achieving high accuracies of 95.33% and 91.67%. Furthermore, the effectiveness in downstream tasks also validates the quality of automatic annotations.
@@ B.2 Dataset Statistics
In this section, we present more detailed statistics about our curated MRGen-DB dataset, including the unannotated image-text pairs from Radiopaedia333radiopaeida.org, as well as the mask-annotated data sourced from various open-source datasets.
Figure 6: Data Statistics of Radiopaedia-MRI. (a) Distribution of slice counts across various modalities in Radiopaedia-MRI; (b) Proportional distribution of slices across different regions in Radiopaedia-MRI.
Data without mask annotations. For the image-text pairs from Radiopaedia-MRI, which are used for training the autoencoder and text-guided generation, we allocate 1% of the data as a test set to evaluate reconstruction and generation performance, maximizing the amount of data available for pretraining. As a result, 202,988 samples are used for training, and the test set consists of 2,051 samples. We conduct a statistical analysis of the distribution of modalities in Radiopaedia-MRI, as presented in Figure 6 (a). The free-text modality labels cover approximately 300 categories, providing a diverse set of MRI modalities that form a crucial foundation for MRGen to learn text-guided generation and expand its mask-conditioned generation capabilities towards modalities originally lacking mask annotations. Furthermore, the distribution of images across different regions in Radiopaedia-MRI is presented in Figure 6 (b).
@@
PROMISE12 [34] Prostate T2-weighed 40 1,137 645 10 240 133
LiQA [36] Liver T1-weighted 24 1,718 1,148 6 467 298
Total / / 778 36,710 14,218 192 7,733 3,643
Table 6: Details of Segmentation-annotated Datasets in MRGen-DB. Here, # Vol. represents the number of 3D Volumes, # Slc. denotes the number of 2D slices, and # Slc. w/ mask indicates the number of 2D slices with mask annotations.
B.3 Synthetic Data Statistics
This section presents the statistics of target-domain training samples synthesized by MRGen across various experimental settings. Concretely, we use mask annotations from the entire source-domain dataset (including both training and test sets) as input conditions to generate target-domain images, forming image-mask training pairs. Exceptions include: (i) for the MSD-Prostate [2] dataset, where images of T2 and ADC modalities have already been registered, we restrict inputs to the source-domain training set to prevent data leakage; and (ii) for dataset pairs with CHAOS-MRI-T1 [26] as the target domain, each source-domain mask is used twice to synthesize both T1-InPhase and T1-OutofPhase data. This setup is consistent across all baselines. Additionally, with our proposed autofilter pipeline, MRGen generates 20 candidate images per mask and selects the top two that meet the predefined thresholds. If no samples satisfy the thresholds, all thresholds will be relaxed by 0.10, and the sample of the highest quality is chosen, ensuring full exploitation of source-domain masks. Otherwise, all low-quality generated samples are discarded to avoid noisy data. Finally, the statistics of target-domain training pairs generated by MRGen for each experimental setting are summarized in Table 7.
Source Dataset Source Modality Target Dataset Target Modality # Slices ($\mathcal{D}_{s}$) # Synthetic Data
CHAOS-MRI T1 CHAOS-MRI T2-SPIR 1,294 433
CHAOS-MRI T2-SPIR CHAOS-MRI T1 607 1,118
MSD-Prostate T2 MSD-Prostate ADC 492 775
MSD-Prostate ADC MSD-Prostate T2 492 745
PanSeg T1 PanSeg T2 18,084 2,160
PanSeg T2 PanSeg T1 15,276 2,215
LiQA T1 CHAOS-MRI T2-SPIR 2,185 2,267
CHAOS-MRI T2-SPIR LiQA T1 607 636
MSD-Prostate ADC PROMISE12 T2 602 742
PROMISE12 T2 MSD-Prostate ADC 1,377 1,077
Table 7: Synthetic Data Statistics. Here, # Slices ($\mathcal{D}_{s}$) denotes the number of source-domain samples under each experimental setting, which serves as input for translation-based baselines. Moreover, # Synthetic Data represents the total volume of data generated by MRGen.
@@ Appendix C Implementation Details
In this section, we will provide a comprehensive explanation of the implementation details discussed in the paper. Concretely, Sec. C.1 describes the preprocessing and augmentation strategies applied to the training data. Sec. C.2 elaborates on the details of the autofilter pipeline. Finally, Sec. C.3 outlines the implementation details of various baselines.
@@ C.1 Preprocessing & Augmentation
Data preprocessing. To ensure consistency across data from various sources and modalities, we apply tailored preprocessing strategies as follows: (i) For data from Radiopaedia-MRI, the images are directly rescaled to the range [0, 1]; (ii) For MR images with mask annotations, intensities are clipped to the 0.5 and 99.5 percentiles and rescaled to [0, 1]. After normalization, all data are subsequently rescaled to [-1, 1] for training various components of MRGen, including the autoencoder, diffusion UNet, and mask condition controller. For training downstream segmentation models, images are rescaled to [0, 255] and saved in ‘.png’ format, followed by the official preprocessing configurations of nnUNet [23] and UMamba [40].
Data augmentation. During autoencoder training, we apply random data augmentations to images with a 20% probability. These augmentations include horizontal flipping, vertical flipping, and rotations of 90, 180, 270. In contrast, no data augmentations are applied during the training of the diffusion UNet and mask condition controller. We do not explicitly adopt data balancing, as we empirically find that it does not lead to significant performance changes. For downstream segmentation models, we adhere to the default data augmentation strategies provided by nnUNet [23] and UMamba [40].
@@ C.2 Autofilter Pipeline
When deploying our proposed data engine, MRGen, to synthesize training data for segmentation models, we adopt the off-the-shelf SAM2-Large [48] model to perform automatic interactive segmentation on generated images, with the mask conditions as spatial prompts. Empirically, we observe that SAM2 consistently segments images based on their contours, guided by the provided spatial prompts. Concretely, it produces high-quality pseudo mask annotations for images with contours closely matching mask conditions, while performing poorly for synthesized images that deviate significantly from mask conditions. This characteristic allows our pipeline to automatically filter out samples faithful to the condition masks and discard erroneous ones, thus ensuring the quality of synthesized image-mask pairs. Here, we elaborate on more implementation details of this automatic filtering pipeline, particularly focusing on the generation of MR images that encompass multiple organs.
Specifically, we begin by defining the following thresholds: confidence threshold ($\tau_{\mathrm{conf}}$), IoU score threshold ($\tau_{\mathrm{IoU}}$), average confidence threshold ($\bar{\tau}_{\mathrm{conf}}$), and average IoU threshold ($\bar{\tau}_{\mathrm{IoU}}$). Both the controlling mask ($\mathcal{M^{\prime}}_{t}$) and the generated image ($\mathcal{I^{\prime}}_{t}$) are fed into SAM2. For each organ mask $\mathcal{M^{\prime}}^{i}_{t}$ in $\mathcal{M^{\prime}}_{t}$, SAM2 will output a segmentation map with a confidence score ($s^{i}_{\mathrm{conf}}$), which is then used to calculate the IoU score ($s^{i}_{\mathrm{IoU}}$) against $\mathcal{M^{\prime}}^{i}_{t}$. For each generated sample ($\mathcal{I^{\prime}}_{t}$), it is regarded to be high-quality and aligned with the mask condition if the following conditions are satisfied: $\{s^{i}_{\mathrm{IoU}}\geq\tau_{\mathrm{IoU}},s^{i}_{\mathrm{conf}}\geq\tau_{\mathrm{conf}}\,|\,\forall i\}$, and $\{\bar{s}_{\mathrm{IoU}}\geq\bar{\tau}_{\mathrm{IoU}}$, $\bar{s}_{\mathrm{conf}}\geq\bar{\tau}_{\mathrm{conf}}\}$. Otherwise, the sample will be discarded.
For each conditional mask, we synthesize 20 image candidates and select the best two that satisfy the predefined thresholds. Across all experiments, the thresholds are set as follows: $\tau_{\mathrm{IoU}}=0.70$, $\tau_{\mathrm{conf}}=0.80$, $\bar{\tau}_{\mathrm{IoU}}=0.80$, and $\bar{\tau}_{\mathrm{conf}}=0.90$.
@@ C.3 Baselines
In this section, we introduce the implementation details of representative baselines and discuss other relevant methods by category. Concretely, we first consider the most related ones, including augmentation-based and translation-based methods.
Augmentation-based methods. These approaches [68, 22, 54, 44, 59, 7] typically rely on mixing multi-domain training data or employing meticulously designed data augmentation strategies. Here, we consider the representative one, DualNorm [68]. Following its official implementation, we apply random non-linear augmentation on each source-domain image, to generate a source-dissimilar training sample, and train the dual-normalization model. All preprocessing steps, network architectures, and training strategies adhere to the official recommendations, with the exception that images are resized to 512 $\times$ 512, consistent with other methods. Notably, we evaluate DualNorm on all slices in the test set, offering a more rigorous evaluation compared to the official code, which only considers slices with segmentation annotations.
Translation-based methods. These methods [52, 27, 46, 28] are commonly inspired by CycleGAN [69]; therefore, we compare with open-source CycleGAN [69], UNSB [27], and MaskGAN [46]. We follow their official implementations and training strategies across all experimental settings. Subsequently, source-domain images are translated into the target domain and paired with the source-domain masks to create paired samples for training downstream segmentation models.
Moreover, we have also explored other approaches leveraging the progress of generative models.
Generation-based methods. Existing medical generation models [60, 14, 15, 9, 55] still struggle with complex and challenging abdominal MRI generation. For instance, MAISI [14] and Med-DDPM [9] are tailored for CT and brain MRI synthesis, respectively. To adapt to our task, we finetune MAISI [14] on our data, as a generation-based baseline.
Additionally, we consider other methods aimed at addressing our focused challenge, i.e., segmenting MR images of underrepresented modalities lacking mask annotations. These include few-shot learning approaches, general-purpose segmentation models, and methods incorporating oracle inputs as performance references. Notably, these approaches, to varying degrees, rely on manually annotated target-domain segmentation masks or external datasets. Thus, they should be regarded as references only rather than fair comparisons with the aforementioned methods and our MRGen.
Few-shot methods. Specifically, we compare with a few-shot nnUNet [23] (pre-trained on source-domain data and finetuned on 5% target-domain manually annotated data), as well as UniVerSeg [5] with its official implementation and checkpoint.
General segmentation models. We adopt the official code and checkpoint of TotalSegmentor-MRI [8], which has been trained on extensive manually annotated data and diverse modalities, as a strong general-purpose segmentation baseline.
Models with oracle inputs. We include SAM2-Large [48] as a reference for interactive semi-automatic segmentation, using randomly perturbed oracle boxes as prompts. To simulate the error introduced by manual intervention, the oracle boxes are randomly shifted at each corner, by up to 8% of the image resolution, following MedSAM [39]. Segmentation results are derived in a slice-by-slice and organ-by-organ manner: For each slice with mask annotations, we simulate box prompts for each annotated organ individually. Finally, we also include nnUNet [23] trained exclusively on the target-domain mask-annotated dataset ($\mathcal{D}_{t}$) as an oracle reference, reflecting the performance upper bound with sufficient annotated data.